art leaves at 10am traveling at 50mph. AT 11:30am Jennifer starts in the same direction at 45mph. when will they be 100 miles apart?
Answer by checkley71(8403)

50X=45(X-1.5)+100
50X=45X-67.5+100
50X-45X=32.5
5X=32.5
X=32.5/5
X=6.5 HOURS THEY WILL BE 100 MILES APART.
50*6.5=45*5+100
325=225+100
325=325